A monopolist has the following information, where Q is the quantity of goods and

P is the price:

Market Demand: P = 300 -Q

Marginal Cost: MC = 2Q

Total Cost: TC = 2 +Q~2


No price discrimination: calculate Q", P*, CS, PS, DWL

Perfect price discrimination: calculate Q", P*, CS, PS, DWL

Second-degree price discrimination: calculate $275, $250, $225, $200 for

each first, second, third, forth of 25 units block.



1
Expert's answer
2021-11-30T10:15:28-0500



1. Under no price-discrimination

"TR= P\times Q = 300Q – Q2; MR = 300 – 2Q"


Set MR = MC; 300 – 2Q = 2Q ; Q = 75.

Hence, Q* = 75


Replace Q*=75 to market demand equation;

"P^* = 300 \u2013 75 = 225."

Hence,"P^* = 225"


"CS =\frac{1}{2} \times (300 – 225)\times 75 = 2812.5"


"PS = \frac{1}{2}\times 150\times 75 + (225-150) \times75 =1125"


"DWL = \frac{1}{2} \times(225-150) \times (100-75) = 937.5"


2. Under perfect price-discrimination

P = MC; 300 – Q = 2Q; Q=100. Hence, Q* = 100


Replace Q* = 100 to market demand equation;

P* = 300 – 100 = 200. Hence, P*=200


CS = 0


PS ="\frac{1}{2} \times 300 \times100 = 15000"


DWL = 0


3. Under second degree price-discrimination:

"TR = 275\times25 + 250\times25 + 225\times25 + 200\times25 = 6,875 + 6,250 + 5,625 + 5,000 = 23,750"


"TC = 2 + 100\times100 = \$10,002"


Profit = 23,750 - 10,002 = 13,748


CS = sum of violet triangle ="\frac{1}{2} \times(300-275) \times 25 \times 4 = 1250"


DWL = 0


PS = Profit = 13,748